OpenStudy (anonymous):

A cylinder labeled X has a height of 12 inches and a radius of 10 inches. Another cylinder labeled Y has a height of 36 inches and a radius of 30 inches. The volume of cylinder Y is _______ times the volume of cylinder X.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (kirbykirby):

\(\text{volume of cylinder} = A_b\times h = \pi r^2 \times h\) So what you should do is find the volume of cylinder X and the volume of cylinder Y, then divide the volume of Y by the volume of X to know by how much volume of Y compares to volume of X

OpenStudy (anonymous):

okay 1 moment

OpenStudy (anonymous):

Is the answer 1:26.986 when rounded 1:27

OpenStudy (kirbykirby):

yes 27 is the right number. In fact, if you leave \(\pi\) written as a symbol itself, rather than an approximation, you get exactly 27: \[\frac{32400\pi}{1200\pi}=\frac{32400}{1200}=27 \]

OpenStudy (anonymous):

So, just 27 is the answer?

OpenStudy (kirbykirby):

yes since the volume for Y is \(32400\pi\) which is 27 times the volume of X (\(1200 \pi\))

OpenStudy (kirbykirby):

It just fits how the question is worded, whereas 1:27 is a ratio and doesn't fit as much with the blank in how the sentence is read

OpenStudy (anonymous):

Ok and thanks

OpenStudy (kirbykirby):

=]